OpenStudy (anonymous):

the school day consists of four block classes, each 90 mins long. Lunch is 30 mins. Passing time between classes and/or lunch is 5 mins. 1) if there is a fire drill at a random time during the day, what is the probability that is begins during lunch? 2) At a random time during a class, your teacher poses the "Question of the Day." If you are 10 mins late for class, what is the probability that you miss hearing the question?

OpenStudy (anonymous):

Total time of school is 5*4 + 90*4 + 30 = 410 mins. lunch probability = 30/410 = 7.32% question probability = 10/30 =33.3%. Assuming at any time in one class session.